# Beyond Instruction Following: Evaluating Rule Following of Large Language Models

### v1

@@ Abstract
Although Large Language Models (LLMs) have demonstrated strong instruction-following ability to be helpful, they are further supposed to be controlled and guided by rules in real-world scenarios to be safe, and accurate in responses. This demands the possession of rule-following capability of LLMs. However, few works have made a clear evaluation of the rule-following capability of LLMs. Previous studies that try to evaluate the rule-following capability of LLMs fail to distinguish the rule-following scenarios from the instruction-following scenarios. Therefore, this paper first makes a clarification of the concept of rule-following, and curates a comprehensive benchmark, RuleBench, to evaluate a diversified range of rule-following abilities. Our experimental results on a variety of LLMs show that they are still limited in following rules. Our further analysis provides insights into the improvements for LLMs toward a better rule-following intelligent agent. The data and code can be found at: https://anonymous.4open.science/r/llm-rule-following-B3E3/
@@
Evaluating Rule Following of Large Language Models
@@ 1 Introduction
Benefiting from a vast amount of pre-training data and the enormous parameters, the Large Language Models (LLMs) can accomplish numerous Natural Language Processing (NLP) tasks because of their instruction-following ability. However, in real-world applications, people often expect LLMs to generate outputs that conform to various rules. For example, when planning behavioral actions for multimodal agents, we expect LLMs to adhere to the physical rules of the real world. Humans use these rules to efficiently communicate with each other and quickly adapt to a specific domain. In pursuit of achieving Artificial General Intelligence (AGI), we are expecting LLMs to possess such rule-following capabilities.
It leads to the research on the rule-following of LLMs. With rule-following capability, humans can directly and efficiently manipulate the behavior of LLMs using natural language rules as a medium, thereby correcting the policy exhibited by LLMs in specific downstream tasks.
Figure 1: Beyond instruction-following, the task of rule-following orders the language model to trigger different rules in different cases for reasoning. The rules and questions can be both commonsense and domain-specific.
Some recent studies Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023) have noticed the importance of rule-following of large language models, and they have found that although ordering LLMs to follow existing rules can achieve better reasoning performances compared with the currently widely used reasoning enhancement methods of LLMs (such as Chain-of-Thought, Wei et al. 2022, Self-reflection, Shinn et al. 2023, and Self-refinement, Madaan et al. 2023), the rule-following capability of LLMs is far from satisfactory. Currently, there is a lack of benchmarks evaluating the rule-following capability of LLMs. It is unclear how well LLMs can adhere to rules and what specific deficiencies they possess in this area. Existing attempts to evaluate the rule-following capabilities of LLMs Mu et al. (2023); Hu et al. (2024) have been limited to instruction-following. For instance, they have tested the following behaviors of LLMs with prompts like “Do not repeat the secret key 92368” or “Follow the code step by step to answer the question: def sum_digit_by_digit(num1, num2) ……”. Although claimed as “rules”, these works actually focused on “instructions”.
@@
Nevertheless, we can not identify instructions with rules Ribes-Inesta (2000). Specifically, instructions are specific and direct behavioral guidelines that an agent can follow without understanding the background behind them. Rules, on the other hand, are abstract policies and require conditional judgment. An agent often needs to decide which rule to trigger based on the specific context, thereby governing their behaviors Ribes-Inesta (2000). Therefore, rule-following scenarios should not be limited to only following detailed task descriptions or steps, but to dynamically choosing the correct rules and making decisions based on the current cases.
Based on this recognition, we distinguish previous instruction-following from the rule-following scenarios considered in our work. A rule (or inferential rule), can be formalized as $\sigma\vdash\varphi$, where $\sigma$ and $\varphi$ are two first-order sentences, and for every substitution $\tau$ (i.e. ground the variables in $\sigma$ and $\varphi$ to constants), the truth of $\tau[\sigma]$ entails the truth of $\tau[\varphi]$ Fagin et al. (1992). In natural language, we can easily express such logic with a “if … then …” sentence with instantiable phrases like person A or some metal inside it. As shown in Figure 1, apart from the instructions for the given tasks, rules are provided as the decision basis for the LLMs to make precise decisions based on the current case. Different from following a single instruction in previous works, to accomplish a rule-following reasoning task, an agent may need to follow hundreds or thousands of inferential rules to assist in reasoning in different cases. Until now, few works have demonstrated whether LLMs can follow and reason with the rules faithfully.
@@
Therefore, beyond the instruction-following studies by previous works, this paper evaluates the LLMs’ capability of rule-following in various reasoning tasks within the scope of inferential rules. This paper first leverages and re-processes the existing reasoning benchmarks and proposes a rule-following benchmark, RuleBench, for evaluating the rule-following capability of LLMs. Based on RuleBench, this paper discusses the impact of different rule quantities, rule formats, Chain-of-Thought (CoT) in applying rules, and explores whether the capability of rule-following still exists in counterfactual scenarios. This paper also analyzes the cases where LLMs fail to follow the rules, categorizing them into Triggering Error and Execution error, which stand for the cases where LLMs fail to trigger the correct rule and LLMs fail to execute the correct rule, respectively. We adopt multiple State-of-The-Art LLMs (§4.1) to accomplish the tasks above. Based on the results, as shown in Figure 2, we categorize the rule-following capabilities of LLMs into 5 dimensions (§4.7), including the capabilities of Triggering Rules, Applying Rules, Executing Rules, Following Formal Rules, and Following Counterfactual Rules. We find that while the closed-source LLMs show dominant performances in the scenario of rule-following, some open-source LLMs, like Llama-3-8B, exhibit competitive performances and have balanced capabilities in all dimensions.
In summary, the major contributions of this paper are as follows:
@@
  • We introduce rule-following as a vital capability of LLMs and distinguish it from the previous labors on instruction-following.
@@
  • We leverage and re-process the existing reasoning benchmarks and propose a rule-following benchmark, RuleBench, for evaluating the rule-following capability of LLMs.
@@
  • We evaluated the capabilities of rule-following of multiple State-of-The-Art LLMs on various tasks and rule settings, and categorized their rule-following abilities into 5 dimensions. Based on the results, we analyze the possible reasons that limit the rule-following capabilities of current LLMs and provide some insights into the improvements for LLMs toward a better rule-following intelligent agent.
@@ 2.1 Rule-enhanced LLM Reasoning
While LLMs have demonstrated remarkable zero-shot reasoning capabilities in many downstream tasks, they still generate outputs that do not conform to logic or human preference. Some research studies have found that compared with the reasoning enhancement methods based on LLMs themselves like Chain-of-Thought Wei et al. (2022), Self-reflection Shinn et al. (2023), and Self-refinement Madaan et al. (2023), providing LLMs with relevant rules with Retrieval-Augmented Generation (RAG) paradigm do better in helping them conduct reasoning in the downstream tasks Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023). However, the rule-following capability of LLMs is far from satisfactory. Few works have comprehensively evaluated whether LLMs can benefit from the provided rules under different scenarios and how LLMs can follow rules better. To make up for this gap, this paper conducted a series of experiments to evaluate the rule-following capabilities of several State-of-The-Art LLMs and provide some insights into how LLMs can follow rules better.
@@ 2.2 LLMs Instruction-following
Instruction-following has been generally considered an important capability of LLMs Zhong et al. (2021); Mishra et al. (2021); Wei et al. (2021); Yin et al. (2023), and some previous works have been done to evaluate the instruction-following capability of LLMs Zhou et al. (2023); Qin et al. (2024). However, only a few works have cast their attention to the question of rule-following. Recent works focused on the rule-following capability of LLMs Mu et al. (2023); Hu et al. (2024) failed to distinguish rule-following from instruction-following. This paper instead proposes the scenario of rule-following and sets up useful baselines for future works.
@@ 3 RuleBench
To construct RuleBench, we have leveraged and re-processed the existing reasoning benchmarks for different rule-following scenarios, including relation extraction, content moderation, commonsense QA, science QA, and judgment prediction. The details of the construction of each benchmark are as follows and the prompts used during constructing RuleBench can be found in Appendix A.

### v2

# Beyond Instruction Following: Evaluating Inferential Rule Following of Large Language Models

@@ Abstract
Although Large Language Models (LLMs) have demonstrated strong instruction-following ability, they are further supposed to be controlled and guided by rules in real-world scenarios to be safe, accurate, and intelligent. This demands the possession of inferential rule-following capability of LLMs. However, few works have made a clear evaluation of the inferential rule-following capability of LLMs. Previous studies that try to evaluate the inferential rule-following capability of LLMs fail to distinguish the inferential rule-following scenarios from the instruction-following scenarios. Therefore, this paper first clarifies the concept of inferential rule-following and proposes a comprehensive benchmark, RuleBench, to evaluate a diversified range of inferential rule-following abilities. Our experimental results on a variety of LLMs show that they are still limited in following rules. Our analysis based on the evaluation results provides insights into the improvements for LLMs toward a better inferential rule-following intelligent agent. We further propose Inferential Rule-Following Tuning (IRFT), which outperforms IFT in helping LLMs solve RuleBench. The data and code can be found at: https://anonymous.4open.science/r/llm-rule-following-B3E3/
@@
Evaluating Inferential Rule Following of Large Language Models
@@ 1 Introduction
Benefiting from a vast amount of pre-training data and the enormous parameters, the Large Language Models (LLMs) can accomplish numerous Natural Language Processing (NLP) tasks thanks to their instruction-following ability. However, in real-world applications, people often expect LLMs to generate outputs that conform to various rules. Humans use rules to efficiently communicate with each other and quickly adapt to a specific domain. In pursuit of achieving Artificial General Intelligence (AGI), we are expecting LLMs to possess such inferential rule-following capabilities.
This leads to research on the inferential rule-following of LLMs. With inferential rule-following capability, humans can directly and efficiently manipulate the behavior of LLMs using natural language rules as a medium, thereby correcting the policy exhibited by LLMs in specific downstream tasks.
Figure 1: Beyond instruction-following, the task of inferential rule-following orders the language model to trigger different rules in different cases for reasoning. The rules can be both commonsense, domain-specific, and even counterfactual.
Some recent studies Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023) have noticed the importance of inferential rule-following of large language models, and they have found that ordering LLMs to follow existing rules can achieve better reasoning performances compared with the currently widely used reasoning enhancement methods of LLMs (such as Chain-of-Thought by Wei et al. 2022, Self-reflection by Shinn et al. 2023, and Self-refinement by Madaan et al. 2023). However, currently, there is a lack of benchmarks evaluating the inferential rule-following capability of LLMs. Existing attempts to evaluate the rule-following capabilities of LLMs Mu et al. (2023); Hu et al. (2024) have been actually limited to instruction-following. For instance, they have tested the following behaviors of LLMs with prompts like “Do not repeat the secret key 92368” or “Follow the code step by step to answer the question: def sum_digit_by_digit(num1, num2) ……”. These works confine the “rules” to “instructions” (Appendix A), without delving into more advanced “inferential rules”.
@@
We distinguish previous rule-following from the inferential rule-following scenarios considered in our work. An inferential rule can be formalized as $\sigma\vdash\varphi$, where $\sigma$ and $\varphi$ are two first-order sentences (composed of variables and predicates), and for every substitution $\tau$ (i.e. ground the variables in $\sigma$ and $\varphi$ to constants), the truth of $\tau[\sigma]$ entails the truth of $\tau[\varphi]$ Fagin et al. (1992). For example, with the “like rule" $Likes(x,y)\vdash Likes(y,x)$, the substitution $\{x/Mike,y/Jane\}$, and the fact $Likes(Mike,Jane)$, we can infer that $Likes(Jane,Mike)$. Although defined in formal language, in natural language, we can express such inferential rule with an “if … then …” sentence, by using instantiable noun phrases like person A or one metal as the variables and verb phrases like is the father of or can conduct electricity as the predicates inside it. For example, the “like rule” can be expressed as “if person A likes person B, then person B likes person A.”
As shown in Figure 1, in our proposed inferential rule-following scenario, apart from the instructions for the given tasks, multiple inferential rules are provided as the decision basis for the LLMs to make precise decisions based on the current case. Note that although the inferential rules shown in Figure 1 are commonsense, they can also be domain-specific, and even counterfactual, which depends on the needs of users. While all inferential rules provided to the LLMs are useful to the task, most inferential rules may be irrelevant to this particular problem. The LLMs need to dynamically trigger and apply the relevant rule (golden rule) based on the current case to conduct reasoning. Until now, few works have demonstrated whether LLMs can follow and reason with the inferential rules faithfully.
Figure 3: The different settings evaluated in RuleBench, including rule quantities, rule forms, Chain-of-Thought (CoT) in applying rules, counterfactual rules, and behavior analysis.
@@
Therefore, beyond the instruction-following studies by previous works, this paper evaluates the LLMs’ capability of inferential rule-following in various reasoning tasks within the scope of inferential rules. This paper proposes a rule-following benchmark, RuleBench, for evaluating the inferential rule-following capability of LLMs under multiple inferential rule-following scenarios, including relation extraction, content moderation, commonsense QA, science QA, and judgment prediction. Based on RuleBench, this paper has evaluated multiple State-of-The-Art LLMs (§4.1), discusses the impact of different rule quantities (§4.2), rule forms (§4.3), Chain-of-Thought (CoT) in applying rules (§4.4), and explores whether the capability of inferential rule-following still exists in counterfactual scenarios (§4.5). This paper also analyzes the cases where LLMs fail to follow the rules (§4.6), categorizing them into Triggering Error and Execution error, which stand for the cases where LLMs fail to trigger the golden rule and LLMs fail to execute the golden rule, respectively. Based on the results, as shown in Figure 2, we categorize the inferential rule-following capabilities of LLMs into 5 dimensions (§4.7), including the capabilities of Triggering Rules, Applying Rules, Executing Rules, Following Formal Rules, and Following Counterfactual Rules. Finally, to further improve the inferential rule-following capabilities of LLMs, we propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases (§4.8). The experimental results show that IRFT significantly outperforms IFT. In summary, the major contributions of this paper are as follows:
@@
  • We introduce inferential rule-following as a vital capability of LLMs and distinguish it from the previous labors on instruction-following.
@@
  • We leverage and re-process the existing reasoning benchmarks and propose an inferential rule-following benchmark, RuleBench, for evaluating the inferential rule-following capability of LLMs.
@@
  • We evaluated the capabilities of inferential rule-following of multiple State-of-The-Art LLMs on various tasks and rule settings, and categorized their inferential rule-following abilities into 5 dimensions. Based on the results, we analyze the possible reasons that limit the inferential rule-following capabilities of current LLMs and provide some insights into the improvements for LLMs toward a better inferential rule-following intelligent agent.
  • We propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases. The experimental results show that IRFT significantly outperforms IFT.
@@ 2.1 Rule-enhanced LLM Reasoning
While LLMs have demonstrated remarkable zero-shot reasoning capabilities in many downstream tasks, they still generate outputs that do not conform to logic or human preference. Some research studies have found that compared with the reasoning enhancement methods based on LLMs themselves like Chain-of-Thought Wei et al. (2022), Self-reflection Shinn et al. (2023), and Self-refinement Madaan et al. (2023), providing LLMs with relevant rules with Retrieval-Augmented Generation (RAG) paradigm do better in helping them conduct reasoning in the downstream tasks Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023). However, the inferential rule-following capability of LLMs is far from satisfactory. Few works have comprehensively evaluated whether LLMs can benefit from the provided rules under different scenarios and how LLMs can follow the rules better. To make up for this gap, this paper conducted a series of experiments to evaluate the inferential rule-following capabilities of several State-of-The-Art LLMs and provide some insights into how LLMs can follow rules better.
@@ 2.2 LLMs Instruction-following
Instruction-following has been generally considered an important capability of LLMs Zhong et al. (2021); Mishra et al. (2021); Wei et al. (2021); Yin et al. (2023), and some previous works have been done to evaluate the instruction-following capability of LLMs Zhou et al. (2023); Qin et al. (2024). However, only a few works have cast their attention to the question of inferential rule-following. Recent works focused on the rule-following capability of LLMs Mu et al. (2023); Hu et al. (2024) confined the rule-following to instruction-following. Instead, this paper proposes the scenario of inferential rule-following and sets up useful baselines for future works.
@@ 3 RuleBench
To construct RuleBench, we have leveraged and re-processed the existing reasoning benchmarks for different inferential rule-following scenarios, including relation extraction (CLUTRR, Sinha et al. 2019), content moderation (SALAD, Li et al. 2024), commonsense QA (DEER, Yang et al. 2022 and ULogic, Wang et al. 2024), mathematics QA (TheoremQa, Chen et al. 2023), and judgment prediction (CAIL2018, Xiao et al. 2018; Zhong et al. 2018). The details of the construction of each benchmark and the prompts used during constructing RuleBench can be found in Appendix B.
Under the scenarios introduced above, As shown in Figure 3, RuleBench involves multiple settings of inferential rule-following, to comprehensively evaluate the LLMs from different perspectives. The settings include rule quantity (i.e. how many rules are provided to the LLMs while only one of them is relevant to the current case), rule form (i.e. which form the rules illustrated in, natural language or formal language), the presence of Chain-of-Thought when applying rules (i.e. directly generate the answer based on the question and rules, or trying verbally apply the rule to the question before answering it), and rule factuality (i.e. whether the conclusion of the rule is factual or counterfactual). RuleBench allows us to analyze the failure cases of inferential rule-following from a behavioral perspective, classifying them into Triggering Error (i.e. LLMs fail to trigger the golden rule) and Execution Error (i.e. LLMs success to trigger the golden rule but fail to execute the golden rule).
4 Evaluation
To comprehensively evaluate the inferential rule-following capabilities of LLMs, based on the proposed RuleBench, this paper has designed 5 main parts of experiments. We evaluate the effects of rule quantity (§4.2), rule form (§4.3), the presence of CoT when applying rules (§4.4), and rule factuality (§4.5). Besides, we analyzed the failure cases of inferential rule-following from a behavioral perspective, classifying them into Triggering Error and Execution Error4.6). Based on these evaluation results, we categorize the inferential rule-following capabilities into 5 dimensions and compare the performances of 8 State-of-The-Art LLMs (§4.7). Finally, to further improve the inferential rule-following capabilities of LLMs, we propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases (§4.8).
4.1 Model Selections
Figure 4: The inferential rule-following performance of LLMs under different rule quantities.
Figure 5: The inferential rule-following performance of LLMs with rules of formal language (FOL) and natural language (NL).
Figure 6: The inferential rule-following performance of LLMs when applying rules with or without using Chain-of-Thought.
4.2 Inferential Rules Are Helpful for the Reasoning of LLMs
To evaluate whether inferential rules are helpful for the reasoning of LLMs, we adopt the following settings to test the LLMs.
  • No Rule. This setting simply prompts the LLMs with the original question and without the inferential rules.
  • Golden Rule. This setting prompts the LLMs with the golden rule (i.e. a relevant rule that should be applied to the question) together with the original question.
  • All Rule. This setting is similar to Few Rule while the number of irrelevant rules increases to 30. This setting simulates a scenario where users prompt the LLMs with all possible inferential rules in the tasks instead of the relevant rules retrieved based on the query.
Besides, we find that by following inferential rules, LLMs have better performance improvements on tasks that require complex reasoning, such as CLUTRR and CAIL2018. On the commonsense reasoning tasks, as the LLMs have parametric knowledge, the performance improvements brought by following inferential rules are relatively slim. Moreover, we find that all LLMs fail to follow the inferential rules in the task of TheoremQA, which illustrates the defect of current LLMs that can not follow complex mathematical or physical rules.
4.3 LLMs Prefer Natural Language Rules than Formal Language Rules
To evaluate whether LLMs can follow formal language rules, we transform the natural language rules of each benchmark into the form of First-Order Logic (FOL) by executing deterministic functions or prompting ChatGPT (Appendix B). Then we compare the reasoning performances of LLMs which are prompted by different forms of inferential rules in both zero-shot All Rule and Few Rule settings.
As shown in Figure 5, in most cases, LLMs conduct reasoning better with natural language rules than formal language rules. This aligns with our intuition that LLMs are mostly pre-trained with natural language and thus the inferential rules expressed with natural language are closer to the pre-trained distributions of LLMs than the inferential rules expressed with formal language. Nevertheless, in most cases, LLMs can follow the formal language rules. This reveals the possibility of learning formal language rules from a symbolic reasoning engine and then using LLMs for neural inference.
Figure 7: The inferential rule-following performance of LLMs when following factual and counterfactual rules.
Figure 8: The failure cases of LLMs inferential rule-following are classified into two categories: Triggering Error and Execution Error, which stand for the cases where LLMs fail to trigger the correct rule and LLMs fail to execute the correct rule, respectively.
4.4 Chain of Thought Is Inadequate for LLMs to Apply Rules
Chain-of-Thought Wei et al. (2022) has been widely verified as a useful prompting technique to help LLMs conduct multi-hop reasoning. To evaluate whether LLMs can use CoT to apply inferential rules in the inferential rule-following scenario, we choose the few-shot Golden Rule and Few Rule settings. We created two demonstrations with CoT and two demonstrations without CoT under such settings for LLMs to conduct In-context Learning.
However, as shown in Figure 6, LLMs with CoT have not exhibited stronger inferential rule-following performances in most cases. This may be attributed to the lack of planning of CoT. CoT conducts straightforward reasoning from the question to the answer with multiple reasoning hops. However, when applying the inferential rules, it involves trying to apply each rule to the current question and thinking about whether to execute this rule. Therefore, plain CoT is inadequate for LLMs to apply the inferential rules. Prompting techniques (e.g. Tree of Thought, Yao et al. 2024) or decoding algorithms (e.g. KCTS, Choi et al. 2023) that involve planning steps are needed for helping LLMs to apply the inferential rules.
4.5 LLMs Struggle to Follow Counterfactual Rules
Although we have verified the effectiveness of the inferential rules, it is still unclear whether LLMs completely follow the given inferential rules or merely use their parametric knowledge. Therefore, we designed the scenario of counterfactual rule-following.
To evaluate whether LLMs can follow counterfactual rules, we construct corresponding counterfactual benchmarks and rule sets of CLUTRR, SALAD, ULogic, and CAIL2018. Specifically, we replace the ground truth of each question and the conclusion of the corresponding rule with a random incorrect answer. So in this counterfactual setting, the LLMs are supposed to generate the “incorrect answer” based on the given counterfactual rules.
As shown in Figure 7, in most cases of both Golden Rule and Few Rule settings, LLMs have significant performance drops when following counterfactual rules, compared with following factual rules. These results indicate that the performance improvements brought by following rules are partly attributed to the parametric knowledge of LLMs, besides following inferential rules.
4.6 Behavioral Analysis of LLMs Following Rules
To understand why LLMs fail to follow the given inferential rules in the reasoning process, we made a behavioral analysis of LLMs in the failure cases of LLMs inferential rule-following. Specifically, we adopt the few-shot Few Rule settings for LLMs to follow the rule-applying demonstrations to apply the given inferential rules to the current question. We ordered the LLMs first to choose an inferential rule to follow and then reason with it. By parsing the output of LLMs we can classify the failure cases of LLMs inferential rule-following into two categories: Triggering Error and Execution Error. Triggering Error indicates that the LLMs choose an irrelevant rule for the current case and therefore lead to an incorrect reasoning result. Execution Error indicates that although LLMs have chosen the correct rule for the current case, they fail to draw the correct conclusion of rule body. To faithfully describe the inferential rule-following behavior of LLMs instead of being affected by the parametric knowledge of LLMs, we run the analysis under the counterfactual settings of the selected benchmarks.
From the results shown in Figure 8, we can tell that when tackling different tasks, LLMs exhibit different behaviors in following rules. While rules have a heavy head for triggering (e.g. in CLUTRR and CAIL2018, the rule head will be a series of relational hops among characters), the LLMs are likely to make Triggering Errors. While the rule head is easy and commonsensical (e.g. in SALAD and ULogic), but the conclusion of the rule body is ambiguous or confused (the counterfactual scenario), the LLMs are likely to make Execution Errors.
To avoid Triggering Errors in the scenario of rule-enhanced reasoning with RAG paradigm (§2.1), the rule retriever plays a crucial role. The Triggering Errors can be eliminated if the rule retriever only retrieved the golden rules. However, existing works often employ simple sparse retrievers such as BM25 Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023), which greatly compromises the inferential rule-following performance of LLMs.
4.7 Rule Following Capabilities of LLMs
To make a comprehensive evaluation of the inferential rule-following capability of the LLMs, we categorize the experimental results in the previous sections into 5 dimensions: Executing Rules, Triggering Rules, Following Formal Rules, Applying Rules, and Following Counterfactual Rules. The details of these dimensions are shown in Appendix C.
As shown in Figure 2, while the closed-source LLMs show dominant performances in the scenario of inferential rule-following, some open-source LLMs, like Llama-3-8B, exhibit competitive performances and have balanced capabilities in all dimensions. Among the closed-source LLMs, gpt-4-turbo is more capable of following formal language rules while gpt-3.5-turbo shows a stronger capability of following counterfactual rules.
Generally, LLMs are not very good at inferential rule-following. This may be attributed to the lack of training in inferential rule-following in the current LLMs. As Instruction Fine-Tuning (IFT) has been a standard step in the pipeline of training LLMs and thus ensures their strong instruction-following capability, in the next section, we propose a fine-tuning method to effectively further improve the inferential rule-following capabilities of LLMs.
Rule Setting Fine-tuned CLUTRR SALAD CAIL2018 No Rule base 22.90 20.12 34.94 IFT 53.82 75.84 80.12 Few Rule base 40.36 88.91 63.25 IRFT 89.50 99.00 90.96 Golden Rule base 55.25 99.67 78.92 IRFT 100.0 99.90 87.95
Table 1: The performances of base and fine-tuned Llama-2-7b-chat on three datasets of RuleBench. IFT stands for instruction fine-tuning, i.e. fine-tuning the LLMs only with the questions and answers. IRFT stands for Inferential Rule-Following Tuning, i.e. train the LLMs to infer the answers with both questions and rules.
4.8 Inferential Rule-Following Tuning
To further improve the inferential rule-following capabilities of LLMs, we propose Inferential inferential rule-following Tuning (IRFT). Compared with IFT, IRFT involves inferential rules as a part of the prompt. The inferential rules can be only the golden rule or the golden rule with a few randomly sampled noise rules. This orders the LLMs to learn to infer the answer not only by the parametric knowledge but also by triggering and executing the golden rule. The tuning objective can be formalized as:
$J_{IRFT}=\mathbb{E}_{\begin{subarray}{c}q,r,a\sim p_{train}\\ r_{1},...,r_{n}\sim U(R)\end{subarray}}-\log p(a|[q;r;r_{1},...,r_{n}])$
Where the $q,r,a$ stands for the question, the golden rule, and the answer from the training set, respectively. $r_{i}\sim U(R)$ stands for randomly sampling $n$ rules from the entire rule sets as the noise rules. Based on the training data in RuleBench, we constructed training data in the settings of No Rule for IFT, and Few Rule ($n=2$) & Golden Rule ($n=0$) for IRFT.
As shown in Table 1, our proposed IRFT further significantly improves the performances of LLMs in the inferential rule-following scenarios and greatly outperforms IFT. This indicates that IRFT can effectively teach the LLMs the capabilities of inferential rule-following.
Although IRFT has shown remarkable performances on RuleBench, beyond using IRFT on specific downstream tasks, we are looking forward to extending IRFT to the pre-training stage of LLMs (like IFT), such that it is possible to enable LLMs to master more basic and generalized inferential rule-following capabilities.
5 Conclusion
In this paper, We introduce inferential rule-following as a vital capability of LLMs and distinguish it from the previous labors on instruction-following. We then construct and propose a new benchmark, RuleBench, for evaluating the inferential rule-following capabilities of LLMs. Based on RuleBench, we conduct a series of experiments to evaluate the inferential rule-following capabilities of 8 State-of-The-Art LLMs from different perspectives. We categorize the inferential rule-following capability in 5 dimensions and provide some insights into improvements for LLMs toward a better inferential rule-following intelligent agent. Finally, we propose the Inferential Rule-Following Tuning (IRFT), which further improves the inferential rule-following capabilities of LLMs.
Limitations
Ethics Statement
Our research aims to evaluate the inferential-inferential rule-following capability of LLMs. To mitigate risks associated with some sensitive content in the benchmark, we restrict access to authorized researchers who adhere to strict ethical guidelines. These measures safeguard research integrity while minimizing potential harm.
Appendix A Instructions vs Rules
Nevertheless, we can not confine rules to instructions, or even identify instructions with rules Ribes-Inesta (2000). Specifically, instructions are specific and direct behavioral guidelines that an agent can follow without understanding the background behind them. Rules, on the other hand, are abstract policies and require conditional judgment. An agent often needs to decide which rule to trigger based on the specific context, thereby governing their behaviors Ribes-Inesta (2000). Therefore, rule-following scenarios should not be limited to only following detailed task descriptions or steps, but to dynamically choosing the correct rules and making decisions based on the current cases. Instructions tell LLMs what to reason, while rules tell LLMs how to reason.
Appendix B Details of Constructing RuleBench
Here are the details of constructing each benchmark in RuleBench. The prompts used in this process are shown in Figure 9,10,11,12,13.